the radius of two circles r1=3 and r2=4 and two circles cut cut orthogonally the area between two circles
 

Grade:12

1 Answers

Nishant Vora IIT Patna
7 years ago
Condition for orthogonality of 2 circles is
d^2 = r_1^2 + r_2 ^2
d^2 = 3^2 + 4 ^2
so d=5

So length of common chord is 4.8 ( by pythagoras )

Thanks